riza was 25 when her first son was born. Today the sum of their ages is 105. How old is riza and her son now??
Answer by rapaljer(4551)

Let x= age of the son now
x+25 = age of Riza now

x + x + 25 = 105
2x + 25 = 105
2x=80
x= 40 Age of son
x+25 = 65 Age of Riza

R^2